Find three consecutive integers such that the square of the sum of the smaller two is 105 more than the square of the largest. I would like to know how to solve because I already have the answer which is 6,7,8.
Found 2 solutions by amoresroy, ewatrrr:
Answer by amoresroy(361)

Find three consecutive integers such that the square of the sum of the smaller two is 105 more than the square of the largest. I would like to know how to solve because I already have the answer which is 6,7,8.
Let x = the smallest integer
x+1 = the next integer
x+2 = the larget integer
(x+x+1)^2 = (x+2)^2 + 105
(2x+1)^2 = (x+2)^2 + 105
4x^2+4x+1 = x^2+4x+4+105
4x cancels out
4x^2-x^2 = 105+4-1
3x^2 = 108
x^2 = 36
x = 6
x+1 = 7
x+2 = 8

Answer by ewatrrr(24785)


Hi,
Let x represent the smallest of the three consecutive integers
then (x+1) and (x+2) would be the next two consecutive integers
Write as we Read
[x + (x+1)]^2 = (x+2)^2 +105
solving for x
[2x + 1)]^2 = (x+2)^2 + 105
4x^2 + 4x + 1 = x^2 + 4x + 4 +105
3x^2 - 108 = 0
x^2 = 108/3
x^2 = 36
x = +- 6
three consecutive integers are 6,7,8 Or -6,-5,-4